OpenStudy (anonymous):

The position vectors of points A, B and C are 8i + 4j – 3k, 6i + 3j - 4k and 7i + 5j – 5k, respectively. Find the area of ∆ ABC

OpenStudy (anonymous):

find AB & AC \[Then area of \triangle=\frac{ 1 }{2 }\left| Vector AB \times Vector AC \right|\]

OpenStudy (anonymous):

Hmm, please solve

OpenStudy (anonymous):

\[Vector AB=P.V of B- p.v. of A=\left( 6i+3j-4k \right)-\left( 8i+4j-3k \right)\] \[=\left( 6-8 \right)i+\left( 3-4 \right)j+\left( -4+3 \right)k=-2i-j-k\] \[similarly Vector AC=\left( 7-8 \right)i+\left( 5-4 \right)j+\left( -5+3 \right)k\] \[=-i+j-2k\]  now you can find the absolute value.

OpenStudy (anonymous):

Wow but no idea of the absolute value?

OpenStudy (anonymous):

\[area=\frac{ 1 }{ 2}\sqrt{\left( 3 \right)^{2}+\left( -5 \right)^{2}+\left( -3 \right)^{2}}=\frac{ 1 }{2 }\sqrt{9+25+9}=\frac{ \sqrt{43} }{ 2 } units.\]

OpenStudy (anonymous):

yw